Q.no.15 mass of block A is 6kg same as that of mass of block B..find v for natural length of spring?Q.no.16mass of each block is 1kg ..first /rear has v :2m/s K:50N find max compression

Grade:11

1 Answers

Kumari Sanjana
6 years ago
Q.16: maximum compression will occur when both the blocks will have equal velocities , according to momentum conservation,m2=mv+mvv=1According to energy conservation1/2*m(2)^2=2*(1/2m(1)^2+ +1/2kx^2x=โˆš0.04=0.2